### v3

@@ Discussion
Human societies are characterized by emergent behavior that cannot be understood by studying just individuals in isolation. Consensus is one of such emergent group capabilities, that is crucial in the development of languages, social norms, and collective decisions. For humans, Dunbar’s number $N_{c}\approx 200$ sets the maximal number of personal relations we can cultivate and thus also the maximal group size in which consensus, intended as the spontaneous emergence of common social norms, can exist. Studying humans and other primates, researchers have identified a power-law scaling connecting the neocortex ratio to the average group size [30], thus proving the link between cognitive capabilities and the development of large societies.
@@
LLMs are attracting a growing interest in the social sciences for their ability to mimic humans, both at the individual and, as recent studies suggest, at the group level. Indeed, like humans, LLM agents show emergent group properties that are not directly coded in their training process. We argue in this paper that consensus is one of these properties, with LLM agents showing striking similarities with primates including humans. As a first result, we showed that all most advanced LLMs are characterized by a majority-following tendency described by a universal function with a single parameter $\beta$, the majority force. Remarkably, this function depends on the specific model but its analytic form is the same describing magnetic spin systems. Different models typically have different $\beta$, with the less sophisticated ones showing smaller $\beta$, i.e., a weaker majority force and thus a more erratic behavior that prevents the reaching of consensus. The majority force depends not only on the LLM, but also on the group size: it tends to be larger in smaller groups. This evidence and the analogy with the Curie-Weiss model allowed us to compute “Dunbar’s number” of each LLM, a size threshold above which a society of agents driven by that LLM is too large to be able to reach consensus. While less sophisticated models show human-like scaling, the most sophisticated LLMs are capable to reach consensus in groups in the thousands, going beyond what human groups can do without rules and institutions.
@@
These results are important due to the relevance of collective behavior and coordination in social contexts. More research is needed to understand other conditions that lead to the emergence of coordination in LLM societies, especially for other types of opinion dynamics with incentives or unequal information access. The ability of LLMs to reach consensus can be beneficial, for example when aiming at coordinating group activities of LLMs where incentives might not be aligned. When there is no information to guide how to behave, LLM societies could spontaneously reach their own social norms, making their behavior predictable by other agents, despite the absence of an intrinsically preferrable choice. However, this also poses threats, as these norms might not be aligned with human values or in situations where coordination threatens the integrity of a system, such as the case of flash crashes among trading bots. Future research in AI anthropology is needed to understand better how this kind of coordination can happen in practical scenarios beyond the idealized situation we studied here. To work towards responsible AI, we need to investigate systemic risks stemming from the collective behavior of LLM agents, where coordination as we showed here is a first example.
@@ Methods
We implement an opinion dynamics process for binary opinions with memoryless LLM agents.
@@
    the process is then iterated until consensus is reached or until the time reaches the maximum preset limit.
@@
LLMs are characterized by a temperature parameter $T$ determining the variance in the sampling of tokens when they response to prompts, such that higher temperatures sample with more variance to the same prompt. In all the simulations reported in the main text we used $T=0.2$. As detailed in the SI, there are no relevant changes when different values of $T$ are used, but a low temperature ensures reliability in the output format.
@@
The Curie-Weiss (CW) Model is arguably the simplest model of a ferromagnet. It describes a system of $N$ spins that can only have two states, either $s_{i}=+1$ (up) or $s_{i}=-1$ (down), coupled with ferromagnetic interactions, i.e., favoring mutual alignment. Each spin interacts with all the others, as the interaction pattern is a fully connected network. The CW model is the mean-field limit of the well-known Ising model. The magnetization $m$, defined as the average of the spin values $m=\langle s_{i}\rangle$ is the equivalent of the average group opinion. The connection between our LLM based opinion simulations and the CW model derives from the transition probability defined by (1). This expression is indeed analogous to the transition probability of Glauber dynamics [42], which allows to simulate the CW model by means of a Markov chain Monte Carlo approach.
@@
The equilibrium value of the magnetization in the CW model can be calculated by means of the self-consistency equation
$m=\tanh(\beta m).$
@@
The solution of this equation depends on the value of inverse temperature $\beta$ (which corresponds, in the opinion dynamics framework, to the majority force). For $\beta<1$ the only solution is $m=0$, while for $\beta>1$ $m=0$ is an unstable solution, while two new solutions $m=\pm m^{*}(\beta)$ appear. The value $\beta_{c}=1$ is a critical point of a second order phase transition. This means that as soon as $\beta>1$, $m^{*}$ grows gradually with $\beta$, tending to $m^{*}=1$ for large values of $\beta$. It is important to remark that in finite systems, the ability to reach consensus depends both on the value of $m^{*}$ and on the fluctuations around this value. In general, order emerges as soon as $\beta>1$, but the system may still not reach full consensus due to statistical fluctuations being too small. For this reason the condition $\beta(N_{c})=1$ is actually an upper bound to the maximal group size where consensus can be reached, since for values of the majority force close to one, the fluctuations may still be not enough for full consensus to be reached.
[…]
Supplementary information Additional analysis and robustness tests are reported in the Supplementary Information
[…]
Acknowledgements We are grateful to Profs. Andrea Baronchelli, Márton Pósfai and Viola Priesemann for interesting discussions. This project was supported by OpenAI with free API credits under its research programme.
@@
Conflict of interest/Competing interests
All authors declare no competing interests.
@@
Author contribution
GDM performed research, all three authors designed and conceptualized research, and all three contributed to manuscript writing.
Appendix A Bias removal
@@
Appendix B Role of opinion names
@@
Appendix C Role of Model temperature
@@ Appendix D Prompt robustness
Appendix E Role of context window length
@@
Appendix F Majority Counting
@@
    You must reply with the opinion supported by the majority.
    The opinion must be reported between square brackets.
We report the result of this analysis for GPT-4 Turbo in Fig. 6 (right). What we observe is a similar adoption probability, that is however much more steep even when $N$ grows. Also in this case detecting the majority gets harder as the system size grows, but, as shown in the inset, its value stays always well above what observed for the opinion dynamics prompt (prompt 1). We can then conclude that the results we obtained can only partially be mapped to a simple counting problem and that the opinion dynamics setting result in a stronger tendency to avoid majority following.

### v4

@@ Discussion
Human societies are characterized by emergent behavior that cannot be understood by studying only individuals in isolation. The ability to coordinate is one of these emergent group properties that is crucial in the development of languages, social norms, and collective decisions. For humans, Dunbar’s number $N_{c}\approx 200$ sets the maximal number of personal relations we can maintain and thus also the maximal group size in which consensus and coordination, intended as the spontaneous emergence of common social decisions, can exist. Studying humans and other primates, researchers have identified a power-law scaling connecting the neocortex ratio to the average group size [18], thus proving the link between cognitive capabilities and the development of large societies.
@@
LLMs are attracting a growing interest in the social sciences for their ability to mimic humans, both at the individual and, as recent studies suggest, at the group level. Like humans, AI agents show emergent group properties that are not directly coded in their training process. We argue that the ability to coordinate and form cohesive groups is one of these properties, with AI agents showing striking similarities with primates including humans. As a first result, we show that all the most advanced LLMs are characterized by a majority-following trend described by a universal function with a single parameter $\beta$, the majority force. Remarkably, this function depends on the specific model but its analytic form is the same describing magnetic spin systems. Different models typically have different $\beta$, with the less sophisticated ones showing smaller $\beta$, i.e., a weaker majority force and thus a more erratic behavior that prevents coordination. The majority force depends not only on the LLM, but also on the group size: it tends to be larger in smaller groups. This evidence and the analogy with the Curie-Weiss model allowed us to compute “Dunbar’s number” of each LLM, a size threshold above which a society of agents driven by that LLM is too large to be cohesive and tends to split. While most models show human-like scaling with cognitive capabilities, some of the most sophisticated LLMs are capable coordinate in groups in the thousands, going beyond what human groups can do without rules and institutions.
@@
These results are important due to the relevance of collective behavior and coordination in social contexts. More research is needed to understand other conditions that lead to the emergence of coordination in AI agent societies, especially for other types of opinion dynamics with incentives, unequal information access, or other kinds of diversity. The ability of LLMs to coordinate can be beneficial, for example when aiming at coordinating group activities of LLMs where no clear preference for any of the options exists. When there is no information to guide how to behave, AI agent societies could spontaneously reach their own social norms, making their behavior predictable by other agents, despite the absence of an intrinsically preferrable choice. However, this also poses threats, as these norms might not be aligned with human values or in situations where coordination threatens the integrity of a system, such as the case of flash crashes among trading bots. Future research with this AI Anthropology perspective is needed to understand better how this kind of coordination can happen in practical scenarios beyond the idealized situation we studied here. To work towards responsible and trustworthy AI, we need to investigate systemic risks stemming from the collective behavior of AI agents, where coordination as we showed here is a first example.
@@ Methods
We implement an opinion dynamics process for binary opinions with memoryless AI agents.
@@
    the process is then iterated until coordination is reached or until the time reaches the maximum preset limit.
@@
LLMs are characterized by a temperature parameter $T$ determining the variance in the sampling of tokens when they respond to prompts, such that higher temperatures sample with more variance to the same prompt. In all the simulations reported in the main text we used $T=0.2$. As detailed in the SI, there are no relevant changes when different values of $T$ are used, but a low temperature ensures reliability in the output format.
@@
The Curie-Weiss (CW) model is arguably the simplest model of a ferromagnet. It describes a system of $N$ spins that can only have two states, either $s_{i}=+1$ (up) or $s_{i}=-1$ (down), coupled with ferromagnetic interactions, i.e., favoring mutual alignment. Each spin interacts with all the others, as the interaction pattern is a fully connected network. The CW model is the mean-field limit of the well-known Ising model. The magnetization $m$, defined as the average of the spin values $m=\langle s_{i}\rangle$ is the equivalent of the average group opinion. The connection between our LLM based opinion simulations and the CW model derives from the transition probability defined by (1). This expression is indeed equal to the transition probability of Glauber dynamics [23], which allows to simulate the CW model by means of a Markov chain Monte Carlo approach.
@@
The equilibrium value of the magnetization in the CW model, which is reached from any initial configuration, is given by the self-consistency equation
$m^{*}=\tanh(\beta m^{*}).$ (2)
@@
The solution of this equation depends on the value of the inverse temperature $\beta$ (which corresponds, in the opinion dynamics framework, to the majority force). For $\beta<1$ the only solution is $m^{*}=0$, while for $\beta>1$, $m^{*}=0$ is an unstable solution, while two new solutions ±$\pm m^{*}(\beta)$ appear. The value $\beta_{c}=1$ is the critical point of a second order phase transition. This means that as soon as $\beta>1$, $m^{*}$ grows gradually with $\beta$, tending to $m^{*}=1$ for large values of $\beta$, approximately following the equation (see SI)
$m^{*}=\sqrt{3(\beta-1)}.$
Critical group size
For $\beta>\beta_{c}$ the equilibrium value of the magnetization is $|m^{*}|>0$, but this does not mean that coordination ($|m|=1$) is reached. The magnetization fluctuates over time around $m^{*}$, the amplitude of fluctuations decreasing with the number of agents $N$. Full coordination is reached when a fluctuation leads from $m^{*}$ to $|m|=1$. Therefore the ability to coordinate crucially depends on the interplay between the value of $m^{*}$ and $N$. If $N$ is not too "large" (in a sense specified below) normal fluctuations are sufficient to rapidly lead to a fully ordered configuration with $|m|=1$. In this case the time to coordinate grows logarithmically with $N$. On the other hand, if $N$ is too "large", coordination can be achieved only when a large and extremely rare fluctuation around $m^{*}$ occurs. The average time for this rare fluctuation grows exponentially with $N$ and hence coordination becomes practically unreachable in reasonable times. Since the equilibrium collective opinion $m^{*}$ is determined by $\beta$, this implies that a crossover value $\beta_{t}(N)$ separates the two regimes.
To get a more precise picture we need to study the stochastic differential equation governing the out of equilibrium evolution of $m$. If we are interested in the behavior around the equilibrium magnetization $m^{*}\approx 1$, we can approximate it with an Ornstein–Uhlenbeck process
$dm=-(m-m^{*})dt+\sqrt{2D}dW(t),$
with
$D\approx\frac{2}{N}\text{e}^{-2\beta}$
More detailed computations are reported in SI. The collective opinion will move randomly around $m^{*}$ with fluctuations of the order of the variance of $m$, that satisfies
$\text{Var}[m]\approx D\approx\frac{2}{N}\text{e}^{-2\beta}$ (3)
This estimate of the typical fluctuations allows to determine the crossover value $\beta_{t}$ separating a regime ($\beta>\beta_{t}$) where fluctuations of the size of the variance lead to coordination, from a regime ($\beta<\beta_{t}$) where a large rare fluctuation is instead necessary. As shown in the SI it holds
$\beta_{t}\approx\frac{1}{2}\log{N}.$ (4)
(10) is reported in Fig. 3. As expected $\beta_{t}>\beta_{c}=1$ for $N$ larger than a few units. Note that for what concerns the time to coordinate there is little difference between systems with $\beta_{c}<\beta<\beta_{t}$ and "disordered" ones with $\beta<\beta_{c}$. In the former case fluctuations occur around a finite value of the magnetization, in the latter around $m^{*}=0$. But in both cases coordination can be achieved only because of an anomalously large rare fluctuation, requiring exponentially large times to develop. Finally, we can estimate the critical group size $N_{c}$ of AI agents as the size for which the majority force reaches $\beta_{t}$
$\beta(N_{c})=\frac{1}{2}\log{N_{c}}.$
Supplementary information
Additional analysis and robustness tests are reported in the Supplementary Information
@@
Acknowledgments
We are grateful to Profs. Andrea Baronchelli, Márton Pósfai and Viola Priesemann for interesting discussions.
Supplementary Information
Bias removal
@@ Acknowledgments
Role of opinion names
@@
Role of Model temperature
@@
Role of context window length
@@
Majority Counting
Model N=51 N=201
gpt-3.5-turbo-0125 0.95 0.97
gpt-4-0613 0.97 0.97
gpt-4-turbo-2024-04-09 0.95 0.95
gpt-4o-2024-05-13 0.95 0.945
claude-2.0 0.94 0.945
claude-3-haiku-20240307 0.95 0.95
claude-3-sonnet-20240229 0.95 0.965
claude-3-opus-20240229 0.985 0.99
claude-3-5-sonnet-20240620 0.945 0.98
meta-llama-3-70b-instruct 0.96 0.955
Table 2: Success rates for majority counting (easy task)
Model N=51 N=201 N=501
gpt-3.5-turbo-0125 0.84 0.85 0.80
gpt-4-0613 0.89 0.845 0.85
gpt-4-turbo-2024-04-09 0.735 0.79 0.75
gpt-4o-2024-05-13 0.825 0.795 0.76
claude-2.0 0.75 0.715 0.715
claude-3-haiku-20240307 0.76 0.715 0.76
claude-3-sonnet-20240229 0.805 0.735 0.74
claude-3-opus-20240229 0.865 0.95 0.925
claude-3-5-sonnet-20240620 0.825 0.90 0.93
meta-llama-3-70b-instruct 0.805 0.775 0.745
Table 3: Success rates for majority counting (hard task)
@@
    Reply with the most frequent letter appearing in the following sequence. The reply must be reported between square brakets.
We used letters $k$ and $z$ and considered lists of different lengths up to $N=501$. We followed two different procedures for generating the lists:
  • Easy task. We first generate a random fraction $p$ between 00 and $1$ and we then generate a list containing a fraction $p$ of letters $k$ and the remaining $z$. We consider $N=51$ and $N=201$;
  • Hard task. We first generate a random fraction $p$ between $0.4$ and $0.6$ and we then generate a list containing a fraction $p$ of letters $k$ and the remaining $z$. In this case most examples will be almost balanced making it harder for the LLM to identify the most frequent letter. For this harder task we also consider longer lists $N=51,201,501$.
The results of this analysis are reported in Tab. 3 and Tab. 3. The results show substantial differences with respect to the prompt we used in our simulations. First, even models that have a very low majority force and are unable to follow the majority in systems as small as $20$, like GPT 3.5 Turbo or Claude 3 Haiku, perform very well in the task. They manage to identify the majority almost always in the easy task and they perform very well also on the harder task, even when $N=501$. For instance GPT 3.5 Turbo on the hard task with $N=500$ performs better than both GPT-4 Turbo and GPT-4o. Moreover, increasing the system size does not lead to rapid degradation of the performance, as in the case of the “Social” prompt. There are models that perform actually better in the longer lists, like Claude 3.5 Sonnet. These results are a strong indication that the effects we observed are derived from the social nature of the prompt and not just from the performance limitations of the LLMs.
Details on LLMs performances and benchmarks
Figure 8: Critical group size vs Benchmarks. We compare the critical group size $N_{c}$ with six different benchmarks measuring the ability of LLMs. $N_{c}$ shows an exponential correlation with the model capabilities.
The performance of LLMs on the MMLU benchmark (and other tasks) has been taken from the official model information and technical reports. When this information was missing and when possible, we integrated it with independent analysis [43]
ChatGPT:
Claude:
Llama:
We show in Fig.  a comparison between six different benchmarks and the critical group size. These are: MMLU, GPQA, HumanEval, Math, MGSM, Elo Score. In all cases we observe an exponential tendency, with the critical group size growing exponentially as the performances of the model increase.
Critical behavior of Curie Weiss Models
We can derive the critical behavior of the CW model expanding the self-consistent equation $m^{*}=\tanh(\beta m^{*})$ for small values of the equilibrium magnetization. This allows to study the system close to the critical point. By expanding the hyperbolic tangent for small values of $\beta m^{*}$
$\tanh(\beta m)=\beta m-\frac{\beta^{3}m^{3}}{3}+\ldots,$
which, inserted into the self-consistent equation, yields
$\ m^{*}\left(1-\beta+\frac{\beta^{3}(m^{*})^{2}}{3}\right)=0$
The value $m^{*}=0$ is always a solution of this equation, while other real solutions exist only if
$11-\beta+\frac{\beta^{3}(m^{*})^{2}}{3}=0\ \to\ \beta>\beta_{c}=$ 1
From this expansion we can also easily obtain the magnetization close to the transition point
$1-\beta+\frac{\beta^{3}(m^{*})^{2}}{3}=0\ \to\ m^{*}=\sqrt{\frac{3(\beta-1)}{\beta^{3}}}$
and since $\beta$ is close to $\beta_{c}=1$ we can further simplify the expression to
$m^{*}=\sqrt{3(\beta-1)}.$
Derivation of $\beta_{t}(N)$
In order to write the stochastic differential equation governing the evolution of the collective opinion $m$, we define the right and left transition probabilities $R(m)$ and $L(m)$. The former gives the probability for the collective opinion to increase by an infinitesimal value $\delta m=2/N$ in a single infinitesimal update $\delta t=1/N$; analogously the latter gives the probability of $m$ to decrease of the same quantity in the same infinitesimal interval. From the adoption probability $P(m)=0.5+0.5\tanh(\beta m)$ we can write
$R(m)=\frac{N_{-}}{N}P(m)=\frac{1-m}{2}\frac{1}{2}\left[1+\tanh(\beta m)\right],$
where the first factor gives the probability of selecting an agent supporting the second opinion and the second one accounts for the probability of this AI agent to transition to the first opinion. Analogously
$L(m)=\frac{1+m}{2}\left[1-P(m)\right]=\frac{1+m}{2}\frac{1}{2}\left[1-\tanh(\beta m)\right].$
Using $R(m)$ and $L(m)$ we can derive the drift $\nu_{m}$ governing the evolution of the average value of the collective opinion
$\frac{d\langle m\rangle}{dt}=\nu_{m}.$
It is easy to show that the drift satisfies
$\nu_{m}=\frac{\delta m}{\delta t}\left[R(m)-L(m)\right]=2P(m)-m-1=\tanh(\beta m)-m.$ (5)
Analogously we can compute $D_{m}$, the diffusion coefficient of the collective opinion, that satisfies
$D_{m}=\frac{1}{2}\frac{(\delta m)^{2}}{\delta t}\left[R(m)+L(m)\right].$
Using the expressions for $R(m)$ and $L(m)$, te last equation gives
$D_{m}=\frac{1}{N}[1-m\tanh(\beta m)].$ (6)
We are interested in studying fluctuations around the equilibrium solution $m^{*}$ that satisfies the self-consistency equation
$m^{*}=\tanh(\beta m^{*})$ (7)
Expanding (5) for small $\delta m=m-m^{*}$ we find
$\nu_{m}\approx-\delta m\left(1-\beta\text{sech}^{2}(\beta m^{*})\right)\approx-(m-m^{*}),$
where the last approximate equality holds already for $\beta=2$ (for which $m^{*}\approx 0.96$).
With regard to the diffusion coefficient, the expansion of the hyperbolic tangent for large values of its argument
$\tanh(x)\approx 1-2\text{e}^{-2x}.$
inserted into (6) yields at leading order
$D_{m}\approx\frac{2}{N}\text{e}^{-2\beta}\equiv D$
Hence the drift is approximately a linearly restoring force, while the diffusion coefficient is approximately constant. The behavior of the collective opinion around the equilibrium value can therefore be described by an Ornstein–Uhlenbeck process
$dm=-(m-m^{*})dt+\sqrt{2D}dW(t),$
where $W(t)$ is a standard Wiener process. The variance of $m$ then is
$\text{Var}[m]\approx\frac{D}{\gamma}=\frac{2}{N}\text{e}^{-2\beta}.$ (8)
The condition for $\beta_{t}$ is
$m^{*}(\beta_{t})+\sqrt{\text{Var}[m](N,\beta_{t})}=1.$ (9)
Expanding the self-consistent equation 7 for $m^{*}\approx 1$ we obtain
$m^{*}\approx 1-2\text{e}^{-2\beta},$
that inserted into (9) yields
$11-2\text{e}^{-2\beta_{t}}+\sqrt{\frac{2}{N}}\text{e}^{-\beta_{t}}=$ 1 .
Solving for $\beta_{t}$ this gives
$\beta_{t}\approx\frac{1}{2}\log{N}.$ (10)
The existence of a phase transition in the CW model can also be derived directly from the time evolution of the magnetization. An ordered state, and eventually a full coordination, can emerge only if the drift is positive for small values of the magnetization, so that the disordered state $m=0$ is unstable. Expanding for small $\beta m$ we get
$\nu_{m}\approx m(\beta-1),$
from which the existence of a phase transition in $\beta_{c}=1$ follows immediately.